# Oxford OX1 3RH, UK

### v2

@@ Abstract
arxiv: 2404.08510
1 Introduction
Neutrinos play a central role in advancing our understanding of physics and addressing fundamental inquiries in contemporary science. The Hyper-Kamiokande experiment Hyper-Kamiokande:2018ofw aims to conduct precise measurements of charge parity violation within the neutrino sector, a phenomenon believed to be closely linked to the observed matter-antimatter asymmetry in the universe. Similarly, the Deep Underground Neutrino Experiment (DUNE) DUNE:2016hlj ; DUNE:2015lol ; DUNE:2016evb ; DUNE:2016rla ; DUNE:2021tad promises the same, in addition to elucidating the neutrino mass ordering. Whether seeking to ascertain Standard Model parameters or probing for exotic phenomena, significant enhancements in both theoretical frameworks and computational simulations of neutrino-nucleus interactions are imperative. As neutrinos interact mainly with nucleons inside the nucleus, the interaction is subject to nuclear effects, namely the nucleon initial state and final-state interactions (FSIs). These effects are difficult to calculate and can alter the final event topology by changing the number of final state pions with respect to the interactions on free nucleons. Notably, the neutrino flux at DUNE yields a comparable proportion of events with and without pions in the final state, highlighting the pressing need for a generator capable of accurately describing both event topologies.
@@ 1 Introduction
This paper is structured as follows: In Sec. 2 and Sec. 3, we review the TKI measurements and genie models respectively. In Sec. 4, we detail the tuning considerations and procedures. Results are summarized in Sec. 5, highlighting how the data-MC discrepancy in the MINERvA-$\pi^{0}$ TKI measurement MINERvA:2020anu is resolved while maintaining good data-MC agreement elsewhere. We conclude in Sec. 6.
2 TKI measurements
@@
TKI is a methodology based on the conservation of momentum in neutrino interactions. In essence, it involves quantifying the imbalance between the observed transverse momentum of the final-state particles and the expected transverse momentum from neutrino interactions with free nucleons Lu:2015hea ; Lu:2015tcr . This “kinematic mismatch” together with its longitudinal and three-dimensional variations Furmanski:2016wqo ; Lu:2019nmf , and the derived asymmetry Cai:2019jzk , has been a crucial set of observables, establishing a pathway to extract valuable information about the participating particles and the underlying nuclear processes. Recent experimental results from neutrino experiments such as T2K T2K:2018rnz ; T2K:2021naz , MINERvA MINERvA:2018hba ; MINERvA:2019ope ; MINERvA:2020anu ; MINERvA:2021csy , and MicroBooNE MicroBooNE:2022emb ; MicroBooNE:2023cmw ; MicroBooNE:2023tzj ; MicroBooNE:2023wzy ; MicroBooNE:2024tmp , as well as electron scattering experiments such as CLAS CLAS:2021neh , highlight the efficacy of TKI.
Figure 1: Schematic illustration of the TKI variables. Diagram taken from Ref. Lu:2015tcr .
@@
In neutrino scattering off a free nucleon, the sum of the transverse components of the final products is expected to be zero, visualized through a back-to-back configuration between the final-state lepton and hadronic system in the plane transverse to the neutrino direction. Hence, in a neutrino interaction with a nucleus, the transverse momentum imbalance, $\delta p_{\textrm{T}}$ Lu:2015tcr , results from intranuclear dynamics, including Fermi motion and FSIs as shown in Fig. 1. The deviation from being back-to-back is quantified by the coplanarity angle $\delta\phi_{\textrm{T}}$ Lu:2015tcr , while the transverse boosting angle, $\delta\alpha_{\textrm{T}}$ Lu:2015tcr , represents the direction of $\delta p_{\textrm{T}}$ within the transverse plane. Furthermore, analyzing the energy and longitudinal momentum budget Furmanski:2016wqo ; Lu:2019nmf enables the conversion of $\delta p_{\textrm{T}}$ to the emulated nucleon momentum, $p_{\textrm{N}}$, providing further insight into the Fermi motion; this conversion amounts to a correction on the order of $\mathcal{O}(20\%)$ Yang:2023dxk . With one-body currents in the absence of FSIs, $\delta\alpha_{\textrm{T}}$ remains uniform (except for second-order effects, such as variations in the center-of-mass energy), given the isotropic nature of the initial nucleon motion. However, as the final products propagate through the nuclear medium, they experience FSIs, thereby disturbing the isotropy and the Fermi motion peak of the $\delta\alpha_{\textrm{T}}$ and $p_{\textrm{N}}$ ($\delta p_{\textrm{T}}$) distributions, respectively. Hence, $\delta p_{\textrm{T}}$ and $p_{\textrm{N}}$ elucidate the Fermi motion details, while $\delta\alpha_{\textrm{T}}$ characterizes the FSI strength—crucial for understanding medium effects in neutrino interactions. A notable advantage of these observables is their minimal dependence on neutrino energy Lu:2015tcr . Moreover, the double TKI variable, $\delta p_{\textrm{TT}}$ Lu:2015hea , is the projection of $\delta\vec{p}_{\textrm{T}}$ along the axis perpendicular to the lepton scattering plane (hence “double”). In addition to its use for extracting neutrino-hydrogen interactions Lu:2015hea ; Hamacher-Baumann:2020ogq , it has also been applied to study nuclear effects in neutrino pion productions MINERvA:2020anu ; T2K:2021naz . Its equivalent in pionless production, $\delta p_{\textrm{Tx}}$, has been proposed and studied together with its orthogonal companion, $\delta p_{\textrm{Ty}}$, in MINERvA MINERvA:2019ope .
@@
This work surveyed four TKI data sets: T2K $0\pi$ T2K:2018rnz , T2K $\pi^{+}$ T2K:2021naz , MINERvA $0\pi$ MINERvA:2018hba ; MINERvA:2019ope , and MINERvA $\pi^{0}$ MINERvA:2020anu measurements. All four measurements require the presence of one CC muon and at least one proton in the final state. While the T2K-$\pi^{+}$ measurement requires exactly one $\pi^{+}$, and MINERvA-$\pi^{0}$ requires at least one $\pi^{0}$, the other two datasets require the absence of any pions. The definitions of the kinematic cuts for these samples are summarised in Table 1.
@@
3 genie model selection
@@
Free nucleon tuning effort in Ref. GENIE:2021zuu has produced a tune, G18_10a_02_11b, with better description of bubble chamber data. This is the starting point of the model we want to tune. The first effect to apply to this model is adding the proper initial state. Among the available genie initial states, the better choice is the spectral-function-like Correlated Fermi gas (SF-CFG) sfcfg-talk ; sfcfg-GitHubCommit ; GENIE:2021npt because it has the most advanced physics and also has the largest number of degrees of freedom that can be adjusted by the tuning procedure. More details on the SF-CFG can be found in Ref. GENIE:2021npt .
Further improvements to the configuration to be tuned include: 1) using $z$-expansion axial-vector form factor Hill:2010yb rather than the dipole form factor of the Valencia model in QE processes Nieves:2004wx ; and 2) replacing the Valencia model Nieves:2011pp in 2p2h processes with the SuSAv2 Model Gonzalez-Jimenez:2014eqa since it covers larger region of the $q^{0}$, $q^{3}$ phase space.
Given its simplicity and interpretability, the INTRANUKE hA FSI model (hA for short) emerges as the FSI candidate for the model to be tuned. The complete list of the model components are given in Table  2. In the genie code, this configuration is identified as G24_20i_00_000 and it is currently the default genie configuration for a number of LAr based experiments. As G24_20i_00_000 is widely used in the paper, we will call it G24-0 for simplicity.
@@
Nuclear state SF-CFG sfcfg-talk ; sfcfg-GitHubCommit ; GENIE:2021npt
QE Valencia Nieves:2004wx
2p2h SuSAv2 Gonzalez-Jimenez:2014eqa
QE $\Delta S=1$ Pais Pais:1971er
QE $\Delta C=1$ Kovalenko Kovalenko:1990zi
Resonance (RES) Berger-Sehgal Berger:2007rq
Shallow/Deep inelastic scattering (SIS/DIS) Bodek-Yang Bodek:2002vp
DIS $\Delta C=1$ Aivazis-Tung-Olness Aivazis:1991fy
Coherent $\pi$ production Berger-Sehgal Berger:2008xs
Hadronization AGKY Yang:2009zx
FSI INTRANUKE hA Andreopoulos:2015wxa
@@
Comparing the genie predictions for G24-0 against TKI datasets, Figs. 3 and 3, shows that the model fails to accurately describe the MINERvA $\pi^{0}$ MINERvA:2020anu TKI measurement.
4 Tuning initial-state and FSI models on TKI data
This study adopts the tuning procedure outlined in Ref. GENIE:2022qrc , utilizing $N_{\textrm{par}}$ model parameters. The objective is to identify the best fit—that is, the optimal set of parameter values—within the parameter space that minimizes the $\chi^{2}$ between model predictions and data. Given the high dimensionality of the parameter space, conducting a brute-force, point-by-point scan on a grid is infeasible. Therefore, a sufficient number of points is randomly sampled. Each point is used to run a full simulation. The simulation output for each data observable bin is then parameterized using Professor Buckley:2009bj with a polynomial of the model parameters of a chosen order, $N_{\textrm{ord}}$. The minimal number of points ($N_{\textrm{s}}$) required for tuning $N_{\textrm{par}}$ parameters with an order $N_{\textrm{ord}}$ polynomial is determined by the combination formula,
@@ 4 Tuning initial-state and FSI models on TKI data
Given that $N_{\textrm{s}}$ increases factorially, caution is essential when augmenting the number of parameters or the polynomial degree. Given that our tuning incorporates all $2$ SF-CFG and $12$ hA parameters (see Sec. 4.2), a degree-$4$ parameterization necessitates over 6,000 generations. Hence, only up to order $4$ polynomials are explored, and it has shown to reproduce the MC predictions well, as shown in Fig. 4.
Figure 4: Fractional difference for the bin-by-bin cross sections between MC truth and the parameterized approximation with order-4 polynomials, both in the Norm-Shape (NS) DAgostini:1993arp ; Hanson:2005mrg space with AllPar. See Table 4 for the definition of AllPar. The residual has a mean of $0.003$ and a standard deviation of $0.016$.
Furthermore, to circumvent Peelle’s Pertinent Puzzle PPP_FNL , the Norm-Shape (NS) transformation prescription DAgostini:1993arp ; Hanson:2005mrg is applied. The extremal point is then found from a minimization of $\chi^{2}$ between this NS-transformed polynomial approximation and NS-transformed data. In minimization, priors, usually based on systematic uncertainty, can be imposed on each parameter to penalize it from getting too far from its default value. The following subsections elaborate on the specific choice of measurement observables and model parameters in this work.
4.1 Data points
The observables for the reported differential cross sections across the four TKI measurements—T2K-$0\pi$, T2K-$\pi^{+}$, MINERvA-$0\pi$, and MINERvA-$\pi^{0}$—are detailed in Table 3. To identify the most sensitive variables, various combinations were evaluated for tuning purposes. A total of 26 combinations were assessed, with the superset comprising $\delta\alpha_{\textrm{T}}$, $\delta p_{\textrm{T}}$, $\delta\phi_{\textrm{T}}$, $p_{\textrm{N}}$, and $\delta p_{\textrm{TT}}$ (see Table 6 in Appendix A for all combinations). Each combination was evaluated such that non-selected variables—including proton momentum and angle ($p_{\text{p}}$ and $\theta_{\text{p}}$vA:2018hba , as well as $\delta p_{\textrm{Tx}}$ and $\delta p_{\textrm{Ty}}$ MINERvA:2019ope in MINERvA-$0\pi$—were used solely for validation; the model is expected to accurately predict these variables post-tuning. When constructing the combinations, it was observed that $\delta\phi_{\textrm{T}}$ is strongly dependent on neutrino beam energy Lu:2015tcr , which suppresses its sensitivity to nuclear effects. Additionally, correlations exist between various observables, notably between $\delta p_{\textrm{T}}$ and $p_{\textrm{N}}$. Therefore, determining the optimal combination is not straightforward; this study employs a criterion based on the $\chi^{2}$ with the complete (tuned plus validation) observable set (Fig. 5).
@@ 4.1 Data points
Figure 5: Change of $\chi^{2}$ calculated for the full (i.e., tuned plus validation) observable set as a function of the tuned combination (cf. Table 6 in Appendix A). The two model parameter sets (AllPar and RedPar, see Table 4 for definitions) are compared and it can be seen that the respective minima happen at Combi-15 and 26.
4.2 Model parameters
As outlined in Sec. 3, we focus exclusively on the currently tunable parameters within the SF-CFG and hA models. Therefore, $14$ parameters (collectively referred to as the AllPar set) are utilized for tuning, as detailed in Table 4. Nominal values and associated uncertainties for the hA model are taken from Table 17.3 in Ref. Andreopoulos:2015wxa .
@@ 4.2 Model parameters
Table 4: Tuneable parameters and their ranges in the SF-CFG (uppermost group) and hA (lower groups, uncertainties from Ref. Andreopoulos:2015wxa ) models. Parameters to be tuned in the two sets are marked with “✓”’s. See later text for definitions of G24-t and G24-c.
In our analysis of the SF-CFG model, the two most relevant parameters are $R_{\textrm{SRC}}$, representing the fraction of the high-momentum tail from short-range correlations (SRC), notably above the Fermi surface, and $E_{\textrm{RM}}^{\textrm{C}}$, marking the commencement of the nuclear removal energy distribution for carbon. A larger $R_{\textrm{SRC}}$ indicates the presence of more energetic initial nucleons, while an increased $E_{\textrm{RM}}^{\textrm{C}}$ implies that greater energy is necessary to liberate a nucleon from the carbon nucleus, so the product particles will possess lower energy. Given the novelty and limited constraints of the spectral-function-like implementation, this study employs relatively relaxed priors for $R_{\textrm{SRC}}$ and $E_{\textrm{RM}}^{\textrm{C}}$, at 0.12$\pm$0.12 and 0.01$\pm$0.005 GeV, respectively.
genie users employ the hA model Andreopoulos:2015wxa for the majority of neutrino analysis, owing to its straightforward reweighting capability. Unlike cascade models such as the hN model Andreopoulos:2015wxa , this model determines the FSI type for each hadron exactly once, without considering further propagation of rescattered hadrons within the nucleus. Its model parameters are explained as follows.
@@
    where $\sigma_{\textrm{hN,tot}}$ is the total hadron-nucleon cross section and $\rho$ is the number density of the nucleons inside the nucleus. Once a hadron is generated inside the nucleus, it travels a distance $\lambda$ before a probabilistic determination of rescattering occurs, inversely related to $\lambda$. If no rescattering happens, the hadron will be propagated for another $\lambda$ and the dice will be thrown again. The cycle repeats until a rescattering happens or the hadron is propagated far enough to escape the nucleus without any interactions. If a rescattering happens, the type of rescattering is evaluated according to their cross section. The relative probabilities of each type are extracted from stored hadron scattering data LADS:1999dyv ; Navon:1983xj ; Carroll:1976hj ; Clough:1974qt ; BAUHOFF1986429 ; Mashnik:2000up ; Ishibashi:1997gbe . After a specific rescattering type is chosen, the rescattered product will be generated. These rescattered products will not undergo further rescattering and will be considered out of the nucleus as final state particles. The MFP values for pions and nucleons can be modified using scaling factors ($S_{\lambda}^{\pi^{\pm}}$, $S_{\lambda}^{\pi^{0}}$, and $S_{\lambda}^{\textrm{N}}$) detailed in Table 4. Both changed pions, $\pi^{+}$ and $\pi^{-}$, are scaled by the same factor, $S_{\lambda}^{\pi^{\pm}}$, due to isospin symmetry. The $\lambda$ for $\pi^{0}$ is calculated from those of charged pions based on isospin symmetry as well and can be adjusted separately by $S_{\lambda}^{\pi^{0}}$. If the factor is larger than $1$, $\lambda$ will increase and hence the total number of steps required for the hadron to escape the nucleus reduces, thereby reducing the average probability of rescattering. Given the high precision with which total hadron-nucleus cross sections are determined LADS:1999dyv ; Navon:1983xj ; Carroll:1976hj ; Clough:1974qt ; BAUHOFF1986429 , strict Gaussian priors (error—$\sigma$, not to be confused with the various cross sections—of $0.2$, same as the systematic uncertainties shown in Table 4) are applied to the MFP scaling factors. Varying $S_{\lambda}^{\pi^{0}}$ by ±$\pm 2.5~{}\sigma$ significantly affects the MINERvA-$\pi^{0}$ $\textrm{d}\sigma/\textrm{d}p_{\textrm{N}}$, as depicted in Fig. 6. Decreasing MFP naturally leads to increased rescattering. Hence, fewer $\pi^{0}$s can escape the nucleus and fewer events will have $\pi^{0}$ in the final states, thereby reducing the cross section.
@@
    The relative probability of each rescattering type can be adjusted by a scale factor. There are four rescattering types for both nucleons and pions available for tuning. They are charge exchange (CEX), inelastic scattering (INEL), absorption (ABS), and pion production (PIPD). The default energy-dependent probability distributions come from hadron data tuning LADS:1999dyv ; Navon:1983xj ; Carroll:1976hj ; Clough:1974qt ; BAUHOFF1986429 and can be adjusted by respective scaling factors, like $S_{\textrm{CEX}}^{\pi}$ for CEX of all pions, listed in Table 4. The relative probabilities of all four rescattering types will be normalized such that they sum up to one. Scaling factors directly multiply the probability of each type, with new relative probabilities subsequently renormalized. More specifically, suppose the initial fraction of absorption, $f_{\textrm{ABS}}$, is $0.5$, which means that the sum of the other fractions is $f_{\textrm{other}}=f_{\textrm{INEL}}+f_{\textrm{CEX}}+f_{\textrm{PIPD}}=1-0.5=0.5$. If $f_{\textrm{ABS}}$ is scaled up with $S_{\textrm{ABS}}=2$, the new absorption fraction is

### v3

# GENIE Collaboration

@@ Abstract
I Introduction
Neutrinos play a central role in advancing our understanding of physics and addressing fundamental inquiries in contemporary science. The Hyper-Kamiokande experiment [1] aims to conduct precise measurements of charge parity violation within the neutrino sector, a phenomenon believed to be closely linked to the observed matter-antimatter asymmetry in the universe. Similarly, the Deep Underground Neutrino Experiment (DUNE) [2, 3, 4, 5, 6] promises the same, in addition to elucidating the neutrino mass ordering. Whether seeking to ascertain Standard Model parameters or probing for exotic phenomena, significant enhancements in both theoretical frameworks and computational simulations of neutrino-nucleus interactions are imperative. As neutrinos interact mainly with nucleons inside the nucleus, the interaction is subject to nuclear effects, namely the nucleon initial state and final-state interactions (FSIs). These effects are difficult to calculate and can alter the final event topology by changing the number of final state pions with respect to the interactions on free nucleons. Notably, the neutrino flux at DUNE yields a comparable proportion of events with and without pions in the final state, highlighting the pressing need for a generator capable of accurately describing both event topologies.
@@ I Introduction
This paper is structured as follows: In Sec. II and Sec. III, we review the TKI measurements and genie models respectively. In Sec. IV, we detail the tuning considerations and procedures. Results are summarized in Sec. V, highlighting how the data-MC discrepancy in the MINERvA-$\pi^{0}$ TKI measurement [14] is resolved while maintaining good data-MC agreement elsewhere. We conclude in Sec. VI.
II TKI measurements
@@
TKI is a methodology based on the conservation of momentum in neutrino interactions. In essence, it involves quantifying the imbalance between the observed transverse momentum of the final-state particles and the expected transverse momentum from neutrino interactions with free nucleons [12, 13]. This “kinematic mismatch” together with its longitudinal and three-dimensional variations [15, 16], and the derived asymmetry [17], has been a crucial set of observables, establishing a pathway to extract valuable information about the participating particles and the underlying nuclear processes. Recent experimental results from neutrino experiments such as T2K [18, 19], MINERvA [20, 21, 14, 22], and MicroBooNE [23, 24, 25, 26, 27], as well as electron scattering experiments such as CLAS [28], highlight the efficacy of TKI.
Figure 1: Schematic illustration of the TKI variables. Diagram taken from Ref. [13].
@@
In neutrino scattering off a free nucleon, the sum of the transverse components of the final products is expected to be zero, visualized through a back-to-back configuration between the final-state lepton and hadronic system in the plane transverse to the neutrino direction. Hence, in a neutrino interaction with a nucleus, the transverse momentum imbalance, $\delta p_{\textrm{T}}$ [13], results from intranuclear dynamics, including Fermi motion and FSIs as shown in Fig. 1. The deviation from being back-to-back is quantified by the coplanarity angle $\delta\phi_{\textrm{T}}$ [13], while the transverse boosting angle, $\delta\alpha_{\textrm{T}}$ [13], represents the direction of $\delta p_{\textrm{T}}$ within the transverse plane. Furthermore, analyzing the energy and longitudinal momentum budget [15, 16] enables the conversion of $\delta p_{\textrm{T}}$ to the emulated (initial) nucleon momentum, $p_{\textrm{N}}$, providing further insight into the Fermi motion; this conversion amounts to a correction on the order of $\mathcal{O}(20\%)$ [29]. With one-body currents in the absence of FSIs, $\delta\alpha_{\textrm{T}}$ remains uniform (except for second-order effects, such as variations in the center-of-mass energy), given the isotropic nature of the initial nucleon motion. However, as the final products propagate through the nuclear medium, they experience FSIs, thereby disturbing the isotropy and the Fermi motion peak of the $\delta\alpha_{\textrm{T}}$ and $p_{\textrm{N}}$ ($\delta p_{\textrm{T}}$) distributions, respectively. Hence, $\delta p_{\textrm{T}}$ and $p_{\textrm{N}}$ elucidate the Fermi motion details, while $\delta\alpha_{\textrm{T}}$ characterizes the FSI strength—crucial for understanding medium effects in neutrino interactions. A notable advantage of these observables is their minimal dependence on neutrino energy [13]. Moreover, the double TKI variable, $\delta p_{\textrm{TT}}$ [12], is the projection of $\delta\vec{p}_{\textrm{T}}$ along the axis perpendicular to the lepton scattering plane (hence “double”). In addition to its use for extracting neutrino-hydrogen interactions [12, 30], it has also been applied to study nuclear effects in neutrino pion productions [14, 19]. Its equivalent in pionless production, $\delta p_{\textrm{Tx}}$, has been proposed and studied together with its orthogonal companion, $\delta p_{\textrm{Ty}}$, in MINERvA [21].
@@
This work surveyed four TKI data sets: T2K $0\pi$ [18], T2K $\pi^{+}$ [19], MINERvA $0\pi$ [20, 21], and MINERvA $\pi^{0}$ [14] measurements. All four measurements require the presence of one CC muon and at least one proton in the final state. While the T2K-$\pi^{+}$ measurement requires exactly one $\pi^{+}$, and MINERvA-$\pi^{0}$ requires at least one $\pi^{0}$, the other two datasets require the absence of any pions. The definitions of the kinematic cuts for these samples are summarized in Table 1.
@@
III genie model selection
@@
Free nucleon tuning effort in Ref. [9] has produced a tune, G18_10a_02_11b, with better description of bubble chamber data. This is the starting point of the model we want to tune. The first effect to apply to this model is adding the proper initial state. Among the available genie initial state models, the best choice is the spectral-function-like Correlated Fermi Gas model (SF-CFG) [38, 39, 8]. SF-CFG is essentially based on the previously implemented Local Fermi Gas (LFG) model in genie with a larger number of adjustable degrees of freedom for tuning and improved physics. It differs from the previous implementation of LFG in two aspects. “Spectral-function-like” refers to the implementation of removal energy as a varying function rather than a fixed value, while “Correlated” highlights the incorporation of the high-momentum tail above the Fermi momentum due to nucleon-nucleon short-range correlations (SRC), as evidenced by electron-scattering data [40]. These two improvements are considered and modeled in the Valencia Model in Ref. [41]. Hence, SF-CFG resembles closer the initial state of the Valencia model than previous LFG implementations. More details on the SF-CFG can be found in Ref. [8]. Note that the original Valencia Model in Ref. [41] has its own method of modeling FSI. However, since FSI is factorized from other models in implementation in genie, this method is not used. Instead, GENIE develops the INTRANUKE hA FSI model (hA for short) for its simplicity and interpretability, which is the FSI candidate for the model to be tuned in this work.
Further improvements to the configuration to be tuned include: 1) using $z$-expansion axial-vector form factor [42] rather than the dipole form factor of the Valencia model in QE processes [41]; and 2) replacing the Valencia model [43] in 2p2h processes with the SuSAv2 Model [44] since it covers larger region of the $q^{0}$, $q^{3}$ phase space.
The complete list of the model components are given in Table  2. In the genie code, this configuration is identified as G24_20i_00_000 and it is currently the default genie configuration for a number of LAr based experiments. As G24_20i_00_000 is widely used in the paper, we will call it G24-0 for simplicity.
Comparing the genie predictions for G24-0 against TKI datasets, Figs. 3 and 3, shows that the model fails to accurately describe the MINERvA $\pi^{0}$ [14] TKI measurement.
@@
Nuclear state SF-CFG [38, 39, 8]
QE Valencia [41]
2p2h SuSAv2 [44]
QE $\Delta S=1$ Pais [45]
QE $\Delta C=1$ Kovalenko [46]
Resonance (RES) Berger-Sehgal [47]
Shallow/Deep inelastic scattering (SIS/DIS) Bodek-Yang [48]
DIS $\Delta C=1$ Aivazis-Tung-Olness [49]
Coherent $\pi$ production Berger-Sehgal [50]
Hadronization AGKY [51]
FSI INTRANUKE hA [52]
@@
IV Tuning initial-state and FSI models on TKI data
This study adopts the tuning procedure outlined in Ref. [11], utilizing $N_{\textrm{par}}$ model parameters. The objective is to identify the best fit—that is, the optimal set of parameter values—within the parameter space that minimizes the $\chi^{2}$ between model predictions and data. Given the high dimensionality of the parameter space, conducting a brute-force, point-by-point scan on a grid is infeasible. Therefore, a sufficient number of points is randomly sampled. Each point is used to run a full simulation. The simulation output for each data observable bin is then parameterized using Professor [53] with a polynomial of the model parameters of a chosen order, $N_{\textrm{ord}}$. The minimal number of points ($N_{\textrm{s}}$) required for tuning $N_{\textrm{par}}$ parameters with an order $N_{\textrm{ord}}$ polynomial is determined by the combination formula,
@@ IV Tuning initial-state and FSI models on TKI data
Given that $N_{\textrm{s}}$ increases factorially, caution is essential when augmenting the number of parameters or the polynomial degree. Given that our tuning incorporates all $2$ SF-CFG and $12$ hA parameters (see Sec. IV.2), a degree-$4$ parameterization necessitates over 6,000 generations. Hence, only up to order $4$ polynomials are explored, and it has shown to reproduce the MC predictions well, as shown in Fig. 4.
Figure 4: Fractional difference for the bin-by-bin cross sections between MC truth and the parameterized approximation with order-4 polynomials, both in the Norm-Shape (NS) [54, 55] space with AllPar. See Table 4 for the definition of AllPar. The residual has a mean of $0.003$ and a standard deviation of $0.073$.
Furthermore, to circumvent Peelle’s Pertinent Puzzle [56, 36], the Norm-Shape (NS) transformation prescription [54, 55] is applied. The extremal point is then found from a minimization of $\chi^{2}$ between this NS-transformed polynomial approximation and NS-transformed data. In minimization, priors, usually based on systematic uncertainty, can be imposed on each parameter to penalize it from getting too far from its default value. The following subsections elaborate on the specific choice of measurement observables and model parameters in this work.
IV.1 Data points
The observables for the reported differential cross sections across the four TKI measurements—T2K-$0\pi$, T2K-$\pi^{+}$, MINERvA-$0\pi$, and MINERvA-$\pi^{0}$—are detailed in Table 3. To identify the most sensitive variables, various combinations were evaluated for tuning purposes. A total of 26 combinations were assessed, with the superset comprising $\delta\alpha_{\textrm{T}}$, $\delta p_{\textrm{T}}$, $\delta\phi_{\textrm{T}}$, $p_{\textrm{N}}$, and $\delta p_{\textrm{TT}}$ (see Table 6 in Appendix .1 for all combinations). Each combination was evaluated such that non-selected variables—including proton momentum and angle ($p_{\text{p}}$ and $\theta_{\text{p}}$[20], as well as $\delta p_{\textrm{Tx}}$ and $\delta p_{\textrm{Ty}}$ [21] in MINERvA-$0\pi$—were used solely for validation; the model is expected to accurately predict these variables post-tuning. When constructing the combinations, it was observed that $\delta\phi_{\textrm{T}}$ is strongly dependent on neutrino beam energy [13], which suppresses its sensitivity to nuclear effects. Additionally, correlations exist between various observables, notably between $\delta p_{\textrm{T}}$ and $p_{\textrm{N}}$. Therefore, determining the optimal combination is not straightforward; this study employs a criterion based on the $\chi^{2}$ with the complete (tuned plus validation) observable set (Fig. 5).
@@ IV.1 Data points
Figure 5: Change of $\chi^{2}$ calculated for the full (i.e., tuned plus validation) observable set as a function of the tuned combination (cf. Table 6 in Appendix .1). The two model parameter sets (AllPar and RedPar, see Table 4 for definitions) are compared and it can be seen that the respective minima happen at Combi-15 and 26.
IV.2 Model parameters
As outlined in Sec. III, we focus exclusively on the currently tunable parameters within the SF-CFG and hA models. Therefore, $14$ parameters (collectively referred to as the AllPar set) are utilized for tuning, as detailed in Table 4. Nominal values and associated uncertainties for the hA model are taken from Table 17.3 in Ref. [52].
@@ IV.2 Model parameters
Table 4: Tuneable parameters and their ranges in the SF-CFG (uppermost group) and hA (lower groups, uncertainties from Ref. [52]) models. Parameters to be tuned in the two sets are marked with “✓”’s. See later text for definitions of G24-t and G24-c.
In our analysis of the SF-CFG model, the two most relevant parameters are $R_{\textrm{SRC}}$, representing the fraction of the high-momentum tail from SRC, notably above the Fermi surface, and $E_{\textrm{RM}}^{\textrm{C}}$, marking the commencement of the nuclear removal energy distribution for carbon. A larger $R_{\textrm{SRC}}$ indicates the presence of more energetic initial nucleons, while an increased $E_{\textrm{RM}}^{\textrm{C}}$ implies that greater energy is necessary to liberate a nucleon from the carbon nucleus, so the product particles will possess lower energy. Given the novelty and limited constraints of the spectral-function-like implementation, this study employs relatively relaxed priors for $R_{\textrm{SRC}}$ and $E_{\textrm{RM}}^{\textrm{C}}$, at 0.12$\pm$0.12 and 0.01$\pm$0.005 GeV, respectively.
genie users employ the hA model [52] for the majority of neutrino analysis, owing to its straightforward reweighting capability. Unlike cascade models such as the hN model [52], this model determines the FSI type for each hadron exactly once, without considering further propagation of rescattered hadrons within the nucleus. Its model parameters are explained as follows.
@@
    where $\sigma_{\textrm{hN,tot}}$ is the total hadron-nucleon cross section and $\rho$ is the number density of the nucleons inside the nucleus. Once a hadron is generated inside the nucleus, it travels a distance $\lambda$ before a probabilistic determination of rescattering occurs, inversely related to $\lambda$. If no rescattering happens, the hadron will be propagated for another $\lambda$ and the dice will be thrown again. The cycle repeats until a rescattering happens or the hadron is propagated far enough to escape the nucleus without any interactions. If a rescattering happens, the type of rescattering is evaluated according to their cross section. The relative probabilities of each type are extracted from stored hadron scattering data [57, 58, 59, 60, 61, 62, 63]. After a specific rescattering type is chosen, the rescattered product will be generated. These rescattered products will not undergo further rescattering and will be considered out of the nucleus as final state particles. The MFP values for pions and nucleons can be modified using scaling factors ($S_{\lambda}^{\pi^{\pm}}$, $S_{\lambda}^{\pi^{0}}$, and $S_{\lambda}^{\textrm{N}}$) detailed in Table 4. Both changed pions, $\pi^{+}$ and $\pi^{-}$, are scaled by the same factor, $S_{\lambda}^{\pi^{\pm}}$, due to isospin symmetry. The $\lambda$ for $\pi^{0}$ is calculated from those of charged pions based on isospin symmetry as well and can be adjusted separately by $S_{\lambda}^{\pi^{0}}$. If the factor is larger than $1$, $\lambda$ will increase and hence the total number of steps required for the hadron to escape the nucleus reduces, thereby reducing the average probability of rescattering. Given the high precision with which total hadron-nucleus cross sections are determined [57, 58, 59, 60, 61], strict Gaussian priors (error—$\sigma$, not to be confused with the various cross sections—of $0.2$, same as the systematic uncertainties shown in Table 4) are applied to the MFP scaling factors. Varying $S_{\lambda}^{\pi^{0}}$ by ±$\pm 2.5~{}\sigma$ significantly affects the MINERvA-$\pi^{0}$ $\textrm{d}\sigma/\textrm{d}p_{\textrm{N}}$, as depicted in Fig. 6. Decreasing MFP naturally leads to increased rescattering. Hence, fewer $\pi^{0}$s can escape the nucleus and fewer events will have $\pi^{0}$ in the final states, thereby reducing the cross section.
@@
    The relative probability of each rescattering type can be adjusted by a scale factor. There are four rescattering types for both nucleons and pions available for tuning. They are charge exchange (CEX), inelastic scattering (INEL), absorption (ABS), and pion production (PIPD). The default energy-dependent probability distributions come from hadron data tuning [57, 58, 59, 60, 61] and can be adjusted by respective scaling factors, like $S_{\textrm{CEX}}^{\pi}$ for CEX of all pions, listed in Table 4. The relative probabilities of all four rescattering types will be normalized such that they sum up to one. Scaling factors directly multiply the probability of each type, with new relative probabilities subsequently renormalized. More specifically, suppose the initial fraction of absorption, $f_{\textrm{ABS}}$, is $0.5$, which means that the sum of the other fractions is $f_{\textrm{other}}=f_{\textrm{INEL}}+f_{\textrm{CEX}}+f_{\textrm{PIPD}}=1-0.5=0.5$. If $f_{\textrm{ABS}}$ is scaled up with $S_{\textrm{ABS}}=2$, the new absorption fraction is